CRY+VIS Interface Meeting
Date : 13:30-14:15
Participants : Takahashi, Kimura, Ikeda, Uchiyama, Hirata, Sato, Aso, Ushiba, Miyoki, Miyo
- IX: Oplev will be finalized in April. Alignment adjustment is scheduled on 4/20. ACC will be installed in April.
Height adjustment (summary).
- IY: Remedy of the pre-isolator will be postponed.
- EX: Height and alignment adjustments are scheduled in 4/21-22.
- EY: Installation of the tower will be restarted in May. Removing of the maraging rod is necessary.
- Corrugate tube fixing jig: They will be prepared by the end of May.
- Commissioning in air: From mid. May for IX.
- Commissioning in vacuum:
- Acceptance check: Meeting was set every Friday to decide the acceptance check items.
- PF actuator: After the commissioning for suspension.
- Jig for FC painting:
- Overall commissioning: Schedule meeting will be held on 4/19.
- Procedure of TM height setting: Prepare the adjustable range tables.
Participants: Takahashi, Fabian, Hirata, Ishizaki, Miyo, Sato, Miyoki, Washimi, Terrence, Aso, Tanaka
- Type-A team: IXV 4/20, EXV 4/21-22
- Type-B team: Remaining work on Type-B/Bp.
IMC suspensions are ready in air  will finish 2011/04/21.
PRs are ready in air  will finish 2011/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2011/11/24.
SRs are ready in air  will finish 2011/12/20.
Output suspensions are ready in air  will finish 2011/12/21.
TMSY is ready in air  will finish 2011/??/??.
TMSX is ready in air  will finish 2011/??/??.
16447 Support for common Tower/Payload model files Type-B(BS) supports TOWER_MASTER and PAYLOAD_MASTER. , satoru.ikeda
16446 Model files have been modified. , satoru.ikeda
16445 Buoyancy effect in PR suspensions. , fabian.arellano
16465 Laser level deviation from the horizontal. , fabian.arellano
16466 Comment , tomotada.akutsu
16468 Tuning of LVDT type accelerometer for ITMX , kouseki.miyo
16365 Wiring the release sensor (outside the vacuum chamber) , naohisa.satou
16438 Comment , naohisa.satou
16421 Height adjustment of ITMX , ryutaro.takahashi
16463 Taking picture of MCo coils , takaaki.yokozawa
16417 Tightening of the inner wire locating clamps. , fabian.arellano
16348 Tightening the locating clamps for the piano wire. , fabian.arellano
16486 Comment , fabian.arellano
This is generated by python script in /users/Miyo/dropbox/Script/scrapping/main.py.
Height adjustment of ITMX (summary).
- Investigation of the stepper motor driver.
- Tuning of the ACCs.
- I checked how large the buoyancy effect is in PR suspensions.
- See klog.
- Some values look very big. I guess this is due to LVDT card replacemet without the adjustment of resistors.
- We checked the accuracy of two laser levels.One of them was very bad and discarded. The other one was very good.
- Miyoki-san has bought another laser level already.
- Using the good laser level and the picomotor we set the pitch of the mirror to -3.3 mrad (Thursday).
- We moved the BF keystone close to the setpoint.
- We released the SF and closed the vacuum chamber.
- The vertical OSEMs were either out of range or close to the end of their ranges. We moved them into range.
- We measured a few IM transfer functions. The system looks healthy.
- Just to check reproducibility, we measured the pitch again and got the expected value of -3.3 mrad (Friday).
- We tested the pitch stability with the Trunp method. We did three rounds and the changes were
- 398 µrad,
- -398 µrad
- -663 µrad.
Then we checked the mirror pitch with respect to the horizontal; we measured 11 mrad. We thought it was too much so we decided to check the accuracy of the laser level.
We repeated the measurement with the good laser level and got 5 mrad, which was more resonable. Using the picomotor we moved it to 3.3 mrad.
- We released the SF and closed the chamber.
- Transfer function measurements showed the system was healthy.
- Although ther suspension was already hanging free, we checked the position of the vertical OSEMs just to make sure the flags were not close to touching. We decided to move V2 and V3 panels.
When we finished we measured the pitch again with the good laser level and didn't get 3.3 mrad, 1.2 mrad! We don't know what could have changed.
We remember another instance in which the pitch changed unexpectedly (after the tightening of the clamps). We need to keep watching it.
- Replacement of a broken LVDT driver (PRM?)
- Set the mirrors to nominal height.
- Terada-san's height marks seem gone.
- PR2 doesn't have one.
- We found one in PR3 chamber which may be Terada-san's one, but there's no indication.
- We haven't checked PR3 chamber yet.
- We must take into account that some LVDT drivers were replaced and the LVDTs were not recalibrated.
- A possible procedure would be something like this:
- Set the SF in nominal height per 3D-CAD.
- Set the laser level at the nominal mirror height.
- Move the BF keystone until the mirror is at the nominal height.
- Take a note of the setpoints and modify medm screens.
- Move vertical OSEMs if necessary.
- Terada-san's height marks seem gone.
- Before pump down set the keystones to heights according to the buoyancy corrections.
- Fishing rods in PR2 and PR3 seem to be working fine.
Suspension commissioning details and plans? [Terrence]
- Has it been discussed? To me, it is still very unclear how it should be done and what people expect. If it has been discussed before, it seems to me it is not transparent.
- I would like to call for a meeting. In the meeting, we should discuss how should SC be done, including but not limited to topics like:
- The scope of the suspension commissioning.
- What are some deliverables?
- How should tasks be done? (E.g. Sensor alignment, noise estimation, etc..)
- What methods should we use? (I propose to use a benchmark so we accept multiple methods, for now.)
- What are some evaluations that should be done? (E.g. We should make a report for each suspension about what has been done and what are some objective quantities that we measured to confirm the performance of the suspensions)
- How to interface with interferometer commissioning? (We shouldn't think of suspension commissioning as something that can be done once and for all. Instead, we need to get input from the interferometer and recursively update our strategies.)
- Someone should roughly plan out the timeline (at least from start to finish).
- I will be drafting a document, which I hope to serve as a handbook/rules for suspension commissioning so everything is clear to everyone.
I found Miyo-san's suspension comissioning document JGW-E2112444-v2. But there are many more things that I would like to address and discuss in detail, for example, how exactly should we do sensor diagonalization? How should we measure performance?
- Which mailing list should I use, to organize the meetings? gokagra?
The week from 4/19
- Takahashi: 4/19-23
- Sato: 4/19-23
- Hirata: 4/19-23
The week from 4/26
- Takahashi: 4/26-28
- Sato: 4/26-5/1
- Hirata: 4/26-28